# The Influence of Clouds and Deuterium-Burning on Brown Dwarf Habitable Zones

### v1

@@ Abstract
To better understand the potential habitability of planets orbiting brown dwarfs, this work presents a new set of equilibrium temperature evolution tracks. Unlike most previous work that relied on analytic scaling relationships for brown dwarf luminosity evolution, we use the outputs of modern brown dwarf evolution models that account for the effects of deuterium burning, cloud formation and dissipation, and the most recent atmospheric opacities. While clouds are present, brown dwarfs cool more slowly than if they did not have clouds, allowing orbiting planets to remain in the habitable zone for millions of years longer than previously estimated. Similarly, we find that during the deuterium-burning phase of brown dwarfs, which also slows the evolution, planets at the same orbital radius but orbiting brown dwarfs of different masses can remain in the habitable zone for the same duration, creating deuterium “sweet spots” for habitability around brown dwarfs near the deuterium-burning limit. For example, at 0.01 AU a planet orbiting both a 0.012 M⊙ and a 0.020 M⊙ brown dwarf stays in the habitable zone for because deuterium burning more strongly affects the cooling of lower-mass brown dwarfs. The size of the effect decreases with decreasing orbital radius, with larger orbital radii having a more pronounced deuterium burning influence. These effects are absent from the analytic cooling approximations used in prior studies of substellar habitable zones and are revealed by our application of modern substellar evolution models.
@@ I Introduction
Most studies of planetary habitability consider planets orbiting main-sequence stars, like Earth orbiting the Sun. However, planets also form and orbit around brown dwarfs (e.g., Zhou et al., 2016; Fontanive et al., 2020; Grieves et al., 2021). Brown dwarfs are substellar objects with masses below the hydrogen burning limit () but above the deuterium burning limit () (Morley et al., 2024). Brown dwarfs occupy the mass range between giant planets and low-mass stars, with sufficient mass to briefly burn deuterium in their cores and effective temperatures spanning K and 250 K (e.g., Burrows et al., 1997; Spiegel et al., 2011; Helling and Casewell, 2014). Unlike main-sequence stars, which are massive enough to fuse hydrogen in their cores, brown dwarfs get cooler and less luminous as they age.
@@
Brown dwarfs are nearly as abundant as main-sequence stars in the universe, making the study of planets around these substellar objects crucial for expanding our understanding of habitability beyond our solar system. This is especially significant considering that M dwarfs constitute the lowest-mass and most abundant class of main-sequence stars in the universe (Kaminski et al., 2025). Observations have revealed that brown dwarf disks exhibit all three early-stage planet formation processes: dust grain growth, crystallization, and settling, indicating that even these very low-mass objects possess the fundamental building blocks for planet formation (Apai et al., 2005), and disks of stellar and sub-stellar mass may be in a fragmentation-limited growth regime in which grain evolution depends on the host star’s mass (Pascucci et al., 2016). Theoretical models predict that brown dwarfs with masses around 0.05 M⊙ could form planets up to 5 M⊕ at orbital distances of 1 AU, though the frequency of planet formation depends critically on disk mass. Earth-mass planets may be relatively common if brown dwarf disks are several Jupiter masses, but extremely rare if typical disks are only a fraction of a Jupiter mass (Payne and Lodato, 2007).
@@
The characteristic that brown dwarfs cool over time, paired with their unique molecular opacity-dominated spectra affects the potential habitability of any planets that orbit them. The thermal emission spectra of brown dwarfs are punctuated by strong molecular features due to water, methane, and other molecules. Thus the atmospheric structure of any habitable planet with an atmosphere will likewise be shaped by the interaction with and absorption of this incident flux. Here, for simplicity, we neglect these spectral features and instead focus on the equilibrium temperature–arising from the integrated incident flux–of planets orbiting brown dwarfs to understand their habitable zone (HZ) lifetimes.
@@
A main sequence star can maintain a planet in the HZ for billions of years, although the HZ moves outwards as the star ages, as shown in Figure 10 (e.g., Kopparapu et al., 2013). This is because such stars get warmer and more luminous over time. On the other hand, a brown dwarf’s HZ move inwards over time as the brown dwarf cools. The cooling proceeds non-uniformly with time, influenced by episodic processes including deuterium burning and the formation or dissipation of condensate clouds, each of which alters the rate of energy loss.
@@
The deuterium burning threshold in brown dwarfs falls approximately in the range of 0.0114 and 0.0144 M⊙ (e.g., Spiegel et al., 2011; Marley et al., 2021; Morley et al., 2024) with the specific limit depending on the details of the atmospheric opacity. A value of 0.013 M⊙ is frequently relied upon as the standard. In the models employed in this study, deuterium burning occurs in brown dwarfs with masses greater than approximately 0.012 to 0.0125 M⊙, for both cloudy and cloudless evolutionary tracks. Brown dwarfs exceeding 0.020 M⊙ will burn through all their deuterium quickly, within 20 million years of their formation, whereas objects below 0.011 M⊙ lack sufficient mass to sustain any appreciable deuterium fusion (e.g., Morley et al., 2019; Saumon et al., 1996; Spiegel et al., 2011). (Morley et al., 2024). Deuterium burning in the mass interval between about 0.012 and 0.020 M⊙ is interesting as it happens more slowly and occurs later in the evolution of brown dwarfs than in more massive objects, making for intriguing effects on HZ lifetimes.
@@
Furthermore, clouds form in brown dwarf atmospheres as they cool, (e.g., Helling and Casewell, 2014; Marley and Robinson, 2015). The atmospheric temperature profiles of brown dwarfs span the condensation points of multiple compounds, resulting in their clouds of various compositions forming and sinking over time as the objects age (e.g., Marley et al., 2002; Metchev et al., 2013). Such clouds alter the atmospheric thermal structure and cooling rates of brown dwarfs as well as their emergent spectra. Optically thick clouds cause the brown dwarf to cool off slower over time than if clouds were not present (e.g., Cooper et al., 2003). Several models have been developed to study cloud structure in brown dwarf atmospheres (e.g., Ackerman and Marley, 2001), providing insights into how clouds influence their thermal structure, spectral properties, and overall evolution (e.g., Saumon and Marley, 2008; Lefèvre et al., 2022; Morley et al., 2024). These studies all document how thermal evolution slows during the time that cloud opacity is most abundant in brown dwarf atmospheres.
@@
The nature and extent of brown dwarf HZ have been studied (e.g., Lingam et al., 2020; Bolmont, 2018; Barnes and Heller, 2013; Andreeshchev and Scalo, 2004) but most of this body of work has only utilized power law fits to brown dwarf evolution model predictions rather than the models themselves to compute equilibrium temperatures for planets orbiting brown dwarfs. For example, Lingam et al. (2020) used the following expression to approximate the effective temperature of brown dwarfs as a function of time:
@@
where tBD is the age of the brown dwarf, MBD is the mass of the brown dwarf, and MJ is the mass of Jupiter. This approximation is from the brown dwarf evolution model of Burrows and Liebert (1993), developed before the first undisputed discovery of a brown dwarf. The power law fit does not capture structure from deuterium burning or cloud formation and dissipation in the evolutionary behavior and produces unrealistically high temperatures at young ages. Differences between this power law model and the models used here are shown in Figure 1.
@@
Similarly, Bolmont (2018) studied the tidal evolution of a planet orbiting a brown dwarf, but their brown dwarf cooling models did not fully account for the effects of deuterium-burning and cloud dynamics in the luminosity evolution of the parent object. Barnes and Heller (2013) used analytical estimates for determining the evolution of brown dwarf luminosity over time to understand its potential habitability and Yates et al. (2017) used analytical estimates to calculate the potential HZ of Y dwarfs.
@@
Here, we aim to refine our understanding of habitable zones around brown dwarfs, and compare these conclusions with the conclusions reached in the above studies. Section II details the key equations and outputs necessary for the analysis. In Section III, we present our findings across four areas: the dependence of planetary equilibrium temperature on orbital distance, the impact on habitable zone lifetime as a function of brown dwarf mass and deuterium-burning, the influence of metallicity on planetary HZ lifetimes, and the overall HZ limit evolution over time. The discussion and conclusions are in sections IV and V, respectively.
@@ II Methods
Because both the Diamondback and Bobcat models do not capture very early brown dwarf evolution () we use the “Red Diamondback” evolution models (Davis et al., 2025) that rely on the SPHINX stellar evolution models (Iyer et al., 2023) for this period of cooling, allowing us to begin evolving all masses at the same age of 1 Myr. At we smoothly transition from Red Diamondback to either the Bobcat or Diamondback tracks, as the case may be. Since clouds do not impact the evolution above 2400 K, the transition to either the Bobcat or Diamondback models is smooth. Overall, using both of those model sets allows for a direct comparison between cloudless and cloudy brown dwarf evolution, with the latter offering a more realistic representation of their true behavior.realistic representation of their true behavior.
@@
where is the effective temperature of the brown dwarf, is the radius of the brown dwarf, is the orbital radius, and is the planet Bond albedo. Throughout this work, ‘habitability’ refers strictly to the radiative habitable zone defined by equilibrium temperature and does not account for atmospheric, geophysical, or tidal processes that may strongly modify surface conditions. The temperature range we consider is 175-275 K, similar to (e.g., Kopparapu et al., 2013). Here we assume a constant Bond albedo of 0.25, but in reality, will vary as the brown dwarf spectrum changes in relation to the planet’s atmospheric absorption profile with time (Marley et al., 1999). A comparison of different Bond albedos is shown in Figure 4. We accept this approximation for now as we are instead focused on the effect of the brown dwarf luminosity evolution. In the future we will consider the impact of the changing shape of the emergent spectra with time.
@@ III Results
As stated in section I, the following section provides an overview of the results across four regimes: equilibrium temperature, habitable zone lifetime as influenced by brown dwarf deuterium burning, habitable zone lifetime as a function of host brown dwarf metallicity, and the evolution of the brown dwarf habitable zone size.
@@ III.1 Equilibrium Temperature
To explore the dependence of equilibrium temperature on orbital radius with the adopted evolution models, we considered three orbital radii: 0.001, 0.01, and 0.1 AU. Figure 3 presents the computed planet equilibrium temperatures as a function of time for the different orbital radii and different brown dwarf evolution model cases. Figure 3 shows that only brown dwarfs below (cloudless evolution) and (cloudy evolution) cool sufficiently for planets at 0.001 AU to enter the habitable zone at ages Gyr. This establishes a clear mass threshold for HZ access at very small orbital radii: more massive brown dwarfs remain too luminous for such planets to ever reach equilibrium temperatures below 275 K within a Hubble time.
@@
III.2 Habitable Zone Lifetime versus Brown Dwarf Mass and Studying Deuterium Burning
In this section, we examine the habitable zone lifetimes of planets orbiting both cloudless and cloudy brown dwarf evolution tracks, considering a range of brown dwarf masses at a range of orbital distances. We identify deuterium-burning “sweet spots” for potential habitability of planets orbiting brown dwarfs. Notably, planets at the same orbital radius around brown dwarfs of different masses near the deuterium-burning limit can remain in the habitable zone for comparable durations. This phenomenon illustrates the significance of deuterium fusion, which temporarily enhances brown dwarf luminosity and extends the window for planetary habitability. We find that brown dwarfs with masses between 0.012 and 0.020 exhibit comparable HZ lifetimes at a fixed orbital distance. This similarity arises because deuterium burning prolongs the cooling timescale of the lower-mass object, an effect that we discuss in greater detail later in this section.
As shown in Figure 4, as the orbital radius increases, the range of brown dwarf masses where planets spend a finite amount of time inside the HZ equilibrium temperature range varies. Around low mass brown dwarfs there is a finite time spent in the HZ at 0.001 AU up to for the clear and for the cloudy evolution, respectively. More massive brown dwarfs never cool off enough for the planet at this orbital radius to enter the HZ temperature range in the age of the Universe. However for the lower brown dwarf masses planets at this orbital radius can spend well in excess of 1 Gyr in the HZ. Varying the Bond albedo from 0.125 to 0.50 demonstrates that planetary albedo directly affects the duration a planet remains in the habitable zone. In the deuterium-burning mass regime, planets with higher albedo spend less time in the habitable zone compared to their lower-albedo counterparts. However, higher-albedo planets can maintain habitable conditions around more massive brown dwarfs, with this effect being modulated by orbital radius.
@@ III.2 Habitable Zone Lifetime versus Brown Dwarf Mass and Studying Deuterium Burning
III.3 The Effects of Metallicity on Brown Dwarf Habitable Zones
Our evolution models allow us to isolate the impact of brown dwarf metallicity on planetary equilibrium temperatures and HZ lifetimes. Figures 5 and 7 illustrate the temperature–age evolution for planets orbiting cloudy evolution brown dwarfs with [M/H] = and at 0.05 AU. Higher-metallicity brown dwarfs cool more slowly and remain warmer at late times, while at very young ages they are slightly cooler than their lower-metallicity counterparts (also illustrated in Figure 1). The behavior at early times reflects differences during the collapse phase: lower metallicity reduces atmospheric opacity, allowing energy to escape more efficiently (e.g., Omukai et al., 2010). At later times, however, the higher opacity of metal-rich atmospheres slows the long-term radiative cooling of the brown dwarf, allowing these objects to remain warmer than lower metallicity brown dwarfs, as they evolve.
The slower long-term cooling of metal-rich brown dwarfs directly translates into extended HZ lifetimes. This trend is shown in Figures 6 and 8, where the duration of habitable conditions increases systematically with metallicity across most brown dwarf masses and orbital radii. As stated previously, clouds increase the temperature of the brown dwarfs which leads to longer planetary habitable lifetimes. A planet located at 0.01 AU around a 0.0125 M⊙ brown dwarf with [M/H] = remains within the HZ for approximately 202 Myr. Similarly, a planet orbiting a 0.020 M⊙ brown dwarf at the same metallicity and distance remains habitable for about 219 Myr. In contrast, at [M/H] = , a planet at 0.01 AU around a 0.0125 M⊙ brown dwarf stays in the HZ for 273 Myr (illustrated as the dotted line in Figure 6), while a similar planet orbiting a 0.020 M⊙ brown dwarf benefits from a significantly longer habitable duration of 333 Myr.
In Figure 8, we see that metallicity significantly influences habitable zone lifetimes in the cloudless evolution models. At 0.01 AU, a planet orbiting a 0.012 M brown dwarf remains in the HZ for approximately 124 Myr at [M/H] = , compared to 158 Myr at [M/H] = . A similar trend is evident for a 0.020 M brown dwarf, where the HZ lifetime increases from 145 Myr at [M/H] = to 192 Myr at [M/H] = .

### v2

@@ Abstract
To better understand the potential habitability of planets orbiting brown dwarfs, this work presents a new set of equilibrium temperature evolution tracks. Unlike most previous work that relied on analytic scaling relationships for brown dwarf luminosity evolution, we use the outputs of modern brown dwarf evolution models that account for the effects of deuterium burning, cloud formation and dissipation, and the most recent atmospheric opacities. While clouds are present, brown dwarfs cool more slowly than if they did not have clouds, allowing orbiting planets to remain in the habitable zone (HZ) for millions of years longer than previously estimated. Similarly, we find that during the deuterium-burning phase of brown dwarfs, which also slows the evolution, planets at the same orbital radius but orbiting brown dwarfs of different masses can remain in the HZ for the same duration, creating deuterium “sweet spots” for habitability around brown dwarfs near the deuterium-burning limit. For example, at 0.01 au a planet orbiting both a 0.012 and a 0.020 M⊙ brown dwarf stays in the HZ for because deuterium-burning more strongly affects the cooling of lower-mass brown dwarfs. The size of the effect decreases with decreasing orbital radius, with larger orbital radii having a more pronounced deuterium burning influence. These effects are absent from the analytic cooling approximations used in prior studies of substellar HZs and are revealed by our application of modern substellar evolution models.
@@ I Introduction
Most studies of planetary habitability consider planets orbiting main-sequence stars, like Earth orbiting the Sun. However, planets also form and orbit around brown dwarfs (e.g., Zhou et al., 2016; Fontanive et al., 2020; Grieves et al., 2021). Brown dwarfs are substellar objects with masses below the hydrogen-burning limit () but above the deuterium-burning limit () (Morley et al., 2024). Brown dwarfs occupy the mass range between giant planets and low-mass stars, with sufficient mass to briefly burn deuterium in their cores and effective temperatures spanning K and 250 K (e.g., Burrows et al., 1997; Spiegel et al., 2011; Helling and Casewell, 2014). Unlike main-sequence stars, which are massive enough to fuse hydrogen in their cores, brown dwarfs get cooler and less luminous as they age.
@@
Brown dwarfs are nearly as abundant as main-sequence stars in the Universe, making the study of planets around these substellar objects crucial for expanding our understanding of habitability beyond our solar system. This is especially significant considering that M dwarfs constitute the lowest-mass and most abundant class of main-sequence stars in the Universe (Kaminski et al., 2025). Observations have revealed that brown dwarf disks exhibit all three early stage planet formation processes: dust grain growth, crystallization, and settling, indicating that even these very low-mass objects possess the fundamental building blocks for planet formation (Apai et al., 2005), and disks of stellar and substellar mass may be in a fragmentation-limited growth regime in which grain evolution depends on the host star’s mass (Pascucci et al., 2016). Theoretical models predict that brown dwarfs with masses around 0.05 M⊙ could form planets up to 5 M⊕ at orbital distances of 1 au, though the frequency of planet formation depends critically on disk mass. Earth-mass planets may be relatively common if brown dwarf disks are several Jupiter masses but extremely rare if typical disks are only a fraction of a Jupiter mass (Payne and Lodato, 2007).
@@
The characteristic that brown dwarfs cool over time, paired with their unique molecular opacity-dominated spectra affects the potential habitability of any planets that orbit them. The thermal emission spectra of brown dwarfs are punctuated by strong molecular features due to water, methane, and other molecules. Thus the atmospheric structure of any habitable planet with an atmosphere will likewise be shaped by the interaction with and absorption of this incident flux. Here, for simplicity, we neglect these spectral features and instead focus on the equilibrium temperature—arising from the integrated incident flux—of planets orbiting brown dwarfs to understand their habitable zone (HZ) lifetimes.
@@
A main-sequence star can maintain a planet in the HZ for billions of years, although the HZ moves outward as the star ages (e.g., Kopparapu et al., 2013). This is because such stars get warmer and more luminous over time. On the other hand, a brown dwarf’s HZ moves inward over time as the brown dwarf cools. The cooling proceeds nonuniformly with time, influenced by episodic processes including deuterium burning and the formation or dissipation of condensate clouds, each of which alters the rate of energy loss.
@@
The deuterium-burning threshold in brown dwarfs falls approximately in the range of 0.0114 and 0.0144 M⊙ (e.g., Spiegel et al., 2011; Marley et al., 2021; Morley et al., 2024) with the specific limit depending on the details of the atmospheric opacity. A value of 0.013 M⊙ is frequently relied upon as the standard. In the models employed in this study, deuterium burning occurs in brown dwarfs with masses greater than approximately 0.012 to 0.0125 M⊙, for both cloudy and cloudless evolutionary tracks. Brown dwarfs exceeding 0.020 M⊙ will burn through all their deuterium quickly, within 20 million years of their formation, whereas objects below 0.011 M⊙ lack sufficient mass to sustain any appreciable deuterium fusion (e.g., Morley et al., 2019; Saumon et al., 1996; Spiegel et al., 2011). (Morley et al., 2024). Deuterium burning in the mass interval between about 0.012 and 0.020 M⊙ is interesting as it happens more slowly and occurs later in the evolution of brown dwarfs than in more massive objects, making for intriguing effects on HZ lifetimes.
@@
Furthermore, clouds form in brown dwarf atmospheres as they cool (e.g., Helling and Casewell, 2014; Marley and Robinson, 2015). The atmospheric temperature profiles of brown dwarfs span the condensation points of multiple compounds, resulting in their clouds of various compositions forming and sinking over time as the objects age (e.g., Marley et al., 2002; Metchev et al., 2013). Such clouds alter the atmospheric thermal structure and cooling rates of brown dwarfs as well as their emergent spectra. Optically thick clouds cause the brown dwarf to cool off slower over time than if clouds were not present (e.g., Cooper et al., 2003). Several models have been developed to study cloud structure in brown dwarf atmospheres (e.g., Ackerman and Marley, 2001), providing insights into how clouds influence their thermal structure, spectral properties, and overall evolution (e.g., Saumon and Marley, 2008; Lefèvre et al., 2022; Morley et al., 2024). These studies all document how thermal evolution slows during the time that cloud opacity is most abundant in brown dwarf atmospheres.
@@
The nature and extent of brown dwarf HZ have been studied (e.g., Lingam et al., 2020; Bolmont, 2018; Barnes and Heller, 2013; Andreeshchev and Scalo, 2004) but most of this body of work has only utilized power-law fits to brown dwarf evolution model predictions rather than the models themselves to compute equilibrium temperatures for planets orbiting brown dwarfs. For example, Lingam et al. (2020) used the following expression to approximate the effective temperature of brown dwarfs as a function of time:
@@
where tBD is the age of the brown dwarf, MBD is the mass of the brown dwarf, and MJ is the mass of Jupiter. This approximation is from the brown dwarf evolution model of Burrows and Liebert (1993), developed before the first undisputed discovery of a brown dwarf. The power-law fit does not capture structure from deuterium burning or cloud formation and dissipation in the evolutionary behavior and produces unrealistically high temperatures at young ages. Differences between this power-law model and the models used here are shown in Figure 1.
@@
Similarly, Bolmont (2018) studied the tidal evolution of a planet orbiting a brown dwarf, but their brown dwarf cooling models did not fully account for the effects of cloud dynamics in the luminosity evolution of the parent object. Barnes and Heller (2013) used models of Baraffe et al. (2003) for the evolution of brown dwarf luminosity over time to understand habitability.
@@
Here, we aim to refine our understanding of HZs around brown dwarfs, and compare these conclusions with the conclusions reached in the above studies. Section II details the key equations and outputs necessary for the analysis. In Section III, we present our findings across four areas: the dependence of planetary equilibrium temperature on orbital distance, the impact on HZ lifetime as a function of brown dwarf mass and deuterium-burning, the influence of metallicity on planetary HZ lifetimes, and the overall HZ limit evolution over time. The discussion and conclusions are in sections IV and V, respectively.
@@ II Methods
Because both the Diamondback and Bobcat models do not capture very early brown dwarf evolution () we use the “Red Diamondback” evolution models (Davis et al., 2025) that rely on the SPHINX stellar evolution models (Iyer et al., 2023) for this period of cooling, allowing us to begin evolving all masses at the same age of 1 Myr. At we smoothly transition from Red Diamondback to either the Bobcat or Diamondback tracks, as the case may be. Since clouds do not impact the evolution above 2400 K, the transition to either the Bobcat or Diamondback models is smooth. Overall, using both of those model sets allows for a direct comparison between cloudless and cloudy brown dwarf evolution, with the latter offering a more realistic representation of their true behavior.
@@
where is the effective temperature of the brown dwarf, is the radius of the brown dwarf, is the orbital radius, and is the planet Bond albedo. Throughout this work, ‘habitability’ refers strictly to the radiative HZ defined by equilibrium temperature and does not account for atmospheric, geophysical, or tidal processes that may strongly modify surface conditions. The temperature range we consider is 175-275 K, similar to (e.g., Kopparapu et al., 2013). Here we assume a constant Bond albedo of 0.25, but in reality, will vary as the brown dwarf spectrum changes in relation to the planet’s atmospheric absorption profile with time (Marley et al., 1999). A comparison of different Bond albedos is presented in the following section. We accept this approximation for now as we are instead focused on the effect of the brown dwarf luminosity evolution. In the future we will consider the impact of the changing shape of the emergent spectra with time.
@@ III Results
As stated in section I, the following section provides an overview of the results across four regimes: equilibrium temperature, HZ lifetime as influenced by brown dwarf deuterium burning, HZ lifetime as a function of host brown dwarf metallicity, and the evolution of the brown dwarf HZ size.
@@ III.1 Equilibrium Temperature
To explore the dependence of equilibrium temperature on orbital radius with the adopted evolution models, we considered three orbital radii: 0.001, 0.01, and 0.1 au. Figure 3 presents the computed planet equilibrium temperatures as a function of time for the different orbital radii and different brown dwarf evolution model cases. Figure 3 shows that only brown dwarfs below (cloudless evolution) and (cloudy evolution) cool sufficiently for planets at 0.001 au to enter the HZ at ages Gyr. This establishes a clear mass threshold for HZ access at very small orbital radii: more massive brown dwarfs remain too luminous for such planets to ever reach equilibrium temperatures below 275 K within a Hubble time.
@@
III.2 HZ Lifetime versus Brown Dwarf Mass and Studying Deuterium Burning
In this section, we examine the HZ lifetimes of planets orbiting both cloudless and cloudy brown dwarf evolution tracks, considering a range of brown dwarf masses at a range of orbital distances. We identify deuterium-burning “sweet spots” for potential habitability of planets orbiting brown dwarfs. Notably, planets at the same orbital radius around brown dwarfs of different masses near the deuterium-burning limit can remain in the HZ for comparable durations. This phenomenon illustrates the significance of deuterium fusion, which temporarily enhances brown dwarf luminosity and extends the window for planetary habitability. We find that brown dwarfs with masses between 0.012 and 0.020 exhibit comparable HZ lifetimes at a fixed orbital distance. This similarity arises because deuterium burning prolongs the cooling timescale of the lower-mass object, an effect that we discuss in greater detail later in this section.
As shown in Figure 4, as the orbital radius increases, the range of brown dwarf masses where planets spend a finite amount of time inside the HZ equilibrium temperature range varies. Around low mass brown dwarfs there is a finite time spent in the HZ at 0.001 au up to for the clear and for the cloudy evolution, respectively. More massive brown dwarfs never cool off enough for the planet at this orbital radius to enter the HZ temperature range in the age of the Universe. However, for the lower brown dwarf masses planets at this orbital radius can spend well in excess of 1 Gyr in the HZ. Varying the Bond albedo from 0.125 to 0.50 demonstrates that planetary albedo directly affects the duration a planet remains in the HZ. In the deuterium-burning mass regime, planets with higher albedo spend less time in the HZ compared to their lower-albedo counterparts. However, higher-albedo planets can maintain habitable conditions around more massive brown dwarfs, with this effect being modulated by orbital radius.
@@ III.2 HZ Lifetime versus Brown Dwarf Mass and Studying Deuterium Burning
III.3 The Effects of Metallicity on Brown Dwarf HZs
Our evolution models allow us to isolate the impact of brown dwarf metallicity on planetary equilibrium temperatures and HZ lifetimes. Figures 5 illustrates the temperature–-age evolution for planets orbiting cloudy evolution brown dwarfs with [M/H] = and at 0.05 au. Higher-metallicity brown dwarfs cool more slowly and remain warmer at late times, while at very young ages they are slightly cooler than their lower-metallicity counterparts (also illustrated in Figure 1). The behavior at early times reflects differences during the collapse phase: lower metallicity reduces atmospheric opacity, allowing energy to escape more efficiently (e.g., Omukai et al., 2010). At later times, however, the higher opacity of metal-rich atmospheres slows the long-term radiative cooling of the brown dwarf, allowing these objects to remain warmer than lower metallicity brown dwarfs, as they evolve.
The slower long-term cooling of metal-rich brown dwarfs directly translates into extended HZ lifetimes. This trend is shown in Figures 6–8. Figure 6 demonstrates that the duration of habitable conditions increases systematically with metallicity across most brown dwarf masses and orbital radii. Similarly, Figure 7 presents the same analysis for the cloudless evolution case, highlighting the reduced habitable lifetimes in the absence of cloud opacity. Figure 8 further illustrates this trend, reinforcing the systematic increase in habitable duration with metallicity. As stated previously, clouds increase the temperature of the brown dwarfs which leads to longer planetary habitable lifetimes. A planet located at 0.01 au around a 0.0125 M⊙ brown dwarf with [M/H] = remains within the HZ for approximately 202 Myr. Similarly, a planet orbiting a 0.020 M⊙ brown dwarf at the same metallicity and distance remains habitable for about 219 Myr. In contrast, at [M/H] = , a planet at 0.01 au around a 0.0125 M⊙ brown dwarf stays in the HZ for 273 Myr (illustrated as the dotted line in Figure 6), while a similar planet orbiting a 0.020 M⊙ brown dwarf benefits from a significantly longer habitable duration of 333 Myr.
In Figure 8, we see that metallicity significantly influences HZ lifetimes in the cloudless evolution models. At 0.01 au, a planet orbiting a 0.012 M brown dwarf remains in the HZ for approximately 124 Myr at [M/H] = , compared to 158 Myr at [M/H] = . A similar trend is evident for a 0.020 M brown dwarf, where the HZ lifetime increases from 145 Myr at [M/H] = to 192 Myr at [M/H] = .